OpenStudy (anonymous):

simplify -14x^3/x^3-5x^4

OpenStudy (anonymous):

x^3/x^3=1

OpenStudy (anonymous):

-14x^3/(x^3-5x^4) ?

OpenStudy (anonymous):

or ( -14x^3/x^3)-5x^4

OpenStudy (anonymous):

-14x^3/(x^3-5x^4) =-14x^3/[x^3*(1-5x)] =-14/(1-5x)

OpenStudy (anonymous):

Ok thats right, can you do this one? x^2-3x-10/x+2

OpenStudy (anonymous):

oh